OpenStudy (anonymous):

Y=2x^+8x+6 in vertex form

OpenStudy (anonymous):

First factor out 2 to rewrite the function as 2(x^2 + 4x + 3). We will now rewrite x^2 + 4x + 3 as x^2 + 4x + 4 - 1 So x^2 + 4x + 3 = (x+2)^2 - 1. Therefore, the final answer is y = 2(x + 2)^2 - 1.